please help me
The length of a field is equal to 120 meters longer than its width.
If the perimeter (distance around) is equal to 560 meters.
Find the dimensions of this field.

Answer by Annabelle1(69)

width=x
length=x+120
Perimeter=x+x+(x+120)+(x+120)
=4x+240
This equals 560
4x+240=560
4x=560-240
4x=320
x=80
Width=80
Length=80+120=200